# Online Combinatorial Optimization with Graphical Dependencies

### v1

@@ 3.3.2 Reducing Game of Googol to $(\frac{1}{2},p)$-sample MRF
Let $\mathcal{I}$ be an instance of the Game of Googol model, and denote by ${\rm OPT}_{GoG}(\mathcal{I})$ and ${\rm ALG}_{GoG}(\mathcal{I})$ our final algorithm’s and the optimal solution cost for $\mathcal{I}$ respectively. Denote by $v^{1}_{1},\dots,v^{1}_{n}$ and $v^{-1}_{1},\dots,v^{-1}_{n}$ the values in $\mathcal{I}$, and by $\bm{\sigma}\sim\mathcal{D}_{\sigma}$ the MRF where $\bm{\sigma}\in\{-1,1\}^{n}$. This defines a sample set $S=(v^{\sigma_{1}}_{1},\dots,v^{\sigma_{n}}_{n})$ and a real set $R=(v^{-\sigma_{1}}_{1},\dots,v^{-\sigma_{n}}_{n})$ in the instance $\mathcal{I}$.
We define the sets $S_{1}=\{v^{1}_{i}\mid\sigma_{i}=1\}$, $R_{1}=\{v^{1}_{i}\mid\sigma_{i}=-1\}$, $S_{2}=\{v^{-1}_{i}\mid\sigma_{i}=-1\}$, and $R_{2}=\{v^{-1}_{i}\mid\sigma_{i}=1\}$. Using these sets, we construct two instances: $\mathcal{I}_{1}=(S_{1},R_{1})$ and $\mathcal{I}_{2}=(S_{2},R_{2})$, where for $i\in\{1,2\}$, we have $S_{i}$ represent the sample set and $R_{i}$ represent the real set (as seen in Figure 2).
[…]
Figure 2: Example of converting a Game of Googol MRF into a $(1/2,p)$-sample MRF instance, for $\bm{\sigma}=(1,1,-1,1,\ldots,-1)$.
For $\mathcal{I}_{1}$, observe that for any value $v_{i}\in S_{1}\cup R_{1}$, we have:
  • $\Pr[v_{i}\in S_{1}]=\frac{1}{2}$ (By the definition of Game of Googol MRF),
  • $\Pr[v_{i}\in S_{1}\mid v_{-i}]\geq\frac{1}{2}\exp\left(-4\Delta\right)$ (By 2.3).
The equivalent property also holds for $\mathcal{I}_{2}$, so both $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are instances for the $(\frac{1}{2},p)$-sample MRF for $p=\frac{1}{2}\exp\left(-4\Delta\right)$. Denote by ${\rm ALG}_{MRF}$ the cost of the $\alpha(p)$-approximation algorithm for the $(\frac{1}{2},p)$-sample MRF model (we ommit the $p$ for brevity) and denote by ${\rm OPT}_{MRF}$ the optimal cost.
@@
[…]
$=2\alpha{\rm OPT}_{GoG}(\mathcal{I})$ Definition of $\text{Definition of }\mathcal{I}_{1},\mathcal{I}_{2}.$
@@ 4 Online Matchings and Combinatorial Auctions
For some distribution $V\sim\mathcal{D}$ and balanced prices $p^{V}_{1},\dots,p^{V}_{n}$, suppose there exist two posted-pricing schemes obtaining expected welfare ${\rm ALG}_{1}$ and ${\rm ALG}_{2}$, respectively, such that
@@
First, notice that Lemma 4.4 ensure that we can always satisfy (1) for any distribution $\mathcal{D}$ that admits $(\alpha,\beta)$-balanced prices. For XOS valuations, Lemma 4.5 gives us $(\alpha,\beta)=(1,1)$-balanced prices that also have a pricing schemes satisfying requirement (2) of Lemma 4.3 with $\gamma=O(\Delta)$ and $\epsilon=1/e$. Similarly, for the hypergraph matching case, Lemma 4.5 gives us $(\alpha,\beta)=(1,k)$-balanced prices that also have a pricing scheme satisfying (2) with $\gamma=O(k^{2}(\Delta+\log k))$ and $\epsilon=\frac{1}{ek}$ which implies the $O((\Delta+\log k)k^{2})$-approximation of the theorem. ∎
@@ 4.1 Tail Contribution
Now, using our MRF conditioning lemma (2.3) and the fact that $R^{V}_{i}$ depends only on $v_{1},\dots,v_{i-1}$, we have ${\mathbb{E}}_{V^{\prime}}[\mathds{1}_{\{j\in O^{V^{\prime}}_{i}\setminus R^{V}_{i}\}}\cdot p^{V^{\prime}}_{j}]\geq e^{-4\Delta}{\mathbb{E}}_{V^{\prime}}[\mathds{1}_{\{j\in O^{V^{\prime}}_{i}\setminus R^{V}_{i}\}}\cdot p^{V^{\prime}}_{j}\mid v_{i}^{\prime}=v_{i}]$. Using this to the above bound and rewriting $V^{\prime}$ as $\hat{V}$ to reflect the conditioning, we have
@@ 4.2.1 XOS Valuations
where each $\mathcal{A}_{i}$ is a subset of $\mathbb{R}_{\geq 0}^{m}$. We show, in Lemma 4.6, how to construct $(1,1)$-balanced prices that also obtain the welfare guarantee of Corollary 4.3.
@@ 4.2.2 Hypergraph Matching
where $M^{V}$ is the max weight hypergraph matching over $E=\{e_{i}:i\in[n]\}$ with weights $w_{e_{i}}=w_{i}$. In our main lemma we give $(1,k)$-balanced prices that also satisfy the requirements of Corollary 4.3, yielding the $O(k^{2}(\Delta+\log k))$ competitive ratio of Theorem 4.1.
@@
To obtain our prices, we first pick $\tau\sim Unif[0,4\Delta+\log k+2]$. Now, for each item $j$, let $\ell_{j}\in\mathbb{Z}$ be the unique integer such that $e^{\tau\ell_{j}}\in[\frac{1}{e^{2}k}b_{j},~{}e^{4\Delta}b_{j})$ for $b_{j}:={\mathbb{E}}_{V}\left[p^{V}_{j}\right]$. Additionally, we sample $X_{\ell}\sim Ber(\frac{1}{k})$ independently for each $\ell\in\mathbb{Z}$ ; this variable indicates whether the item’s price will be set to high or low.
@@
$p_{j}:=\begin{cases}\frac{1}{e}\cdot e^{4\Delta}b_{j}&X_{\ell_{j}}=0\text{ (``high price''}),\\ \frac{1}{e}\cdot e^{\tau\ell_{j}}&X_{\ell_{j}}=1\text{ (``low price'')}.\end{cases}$
First, we show that items which sell frequently at a high price achieve a high enough revenue to account for all their neighboring edges. Let $H^{\tau}:=\left\{j:\Pr[j\text{ sells}\mid X_{\ell_{j}}=0,~{}\tau]\geq\frac{1}{k^{2}}e^{-4\Delta-2}\right\}$ and denote by $\delta(H^{\tau})$ the neighboring edges of the items in the set $H^{\tau}$.
@@
For fixed $\tau$, we have ${\mathbb{E}}\big{[}\sum_{j\in H^{\tau}}\mathds{1}_{\{j\text{ sells}\}}\cdot p_{j}\mid\tau\big{]}\geq\frac{1}{2e^{3}k^{2}}\sum_{\begin{subarray}{c}e_{i}\in\delta(H^{\tau})\end{subarray}}\mathds{1}_{\{e_{i}\in M^{V}\}}\cdot w_{i}$
Proof of Claim 4.8.
For $j\in H^{\tau}$, we have
${\mathbb{E}}[\mathds{1}_{\{j\text{ sells}\}}\cdot p_{j}\mid\tau]$ $\geq{\mathbb{E}}\left[\mathds{1}_{\{j\text{ sells}\}}\cdot e^{4\Delta-1}b_{j}\mid X_{\ell_{j}}=0,~{}\tau\right]\Pr[X_{\ell_{j}}=0\mid\tau]$
@@
Using $b_{j}={\mathbb{E}}\left[\sum_{e_{i}\in\delta(j)}\mathds{1}_{\{e_{i}\in M^{V}\}}w_{i}\right]$ and summing over $j\in H^{\tau}$, we get our claim. ∎
Hence, it suffices to focus on the contribution from buyers $i$ such that $e_{i}\not\in\delta(H^{\tau})$. For such a buyer $i$, define the $\xi_{i}$ to be the event that
@@
In other words, $\xi_{i}$ is the event that the item $j\in e_{i}$ with large $b_{j}$ are all “low-priced,” while all other items are “high-priced.” Let $B^{i}_{low}=\{j\in e_{i}:X_{\ell_{j}}=1\}$ and $B^{i}_{high}=\{j\in e_{i}:X_{\ell_{j}}=0\}$ denote these low and high priced items respectively. Notice that even in the event $\xi_{i}$, all items in $e_{i}$ have price at most $\frac{1}{ek}e^{\tau\ell^{i}_{\max}}$. Additionally, we note that $\xi_{i}$ is independent of the valuation $V$, and
@@
$\geq{\mathbb{E}}_{V}\Bigg{[}\mathds{1}_{\{\xi_{i}\}}\mathds{1}_{\{e_{i}\subseteq R^{V}_{i}\}}\cdot\Big{(}w_{i}-\frac{1}{e}e^{\tau\ell^{i}_{\max}}\Big{)}^{+}\Bigg{]}$
@@
Here, $\chi_{i}$ denotes the event that $\frac{e^{\tau\ell^{i}_{\max}}}{\overline{w}_{i}}\in[\frac{1}{e},1]$. We show the following claim on the term that corresponds to the high-priced items.
@@
${\mathbb{E}}\Big{[}\sum_{j\in[m]}\mathds{1}_{\{j\text{ sold}\}}\cdot p_{j}\Big{]}$ $\geq{\mathbb{E}}\Big{[}\sum_{j}\mathds{1}_{\{X_{\ell_{j}}=1\}}\mathds{1}_{\{j\text{ sold}\}}\cdot\frac{1}{ek}e^{\tau\ell_{j}}\Big{]}$
$={\mathbb{E}}\Big{[}\sum_{e_{i}\in M^{V}}\sum_{j\in e_{i}}\mathds{1}_{\{X_{\ell_{j}}=1\}}\mathds{1}_{\{j\text{ sold}\}}\cdot\frac{1}{ek}e^{\tau\ell_{j}}\Big{]}$
$={\mathbb{E}}\Big{[}\sum_{e_{i}\in M^{V}}\sum_{j\in B^{i}_{low}}\mathds{1}_{\{j\text{ sold}\}}\cdot\frac{1}{ek}e^{\tau\ell_{j}}\Big{]}$
$\geq{\mathbb{E}}\Big{[}\sum_{e_{i}\in M^{V}}\sum_{j\in B^{i}_{low}}\mathds{1}_{\{\xi_{i}\}}\mathds{1}_{\{j\text{ sold}\}}\cdot\frac{1}{ek}e^{\tau\ell^{i}_{\max}}\Big{]}$
$\geq{\mathbb{E}}\Big{[}\sum_{e_{i}\in M^{V}}\mathds{1}_{\{\xi_{i}\}}\mathds{1}_{\{B^{i}_{low}\not\subseteq R^{V}_{i}\}}\cdot\frac{1}{ek}e^{\tau\ell^{i}_{\max}}\Big{]}$
$\geq{\mathbb{E}}\Big{[}\sum_{e_{i}\in M^{V}}\mathds{1}_{\{\chi_{i}\}}\mathds{1}_{\{\xi_{i}\}}\mathds{1}_{\{B^{i}_{low}\not\subseteq R^{V}_{i}\}}\cdot\frac{1}{e^{2}k}\overline{w}_{i}\Big{]}.$
@@
${\mathbb{E}}\Big{[}\sum_{i\not\in\delta(H^{\tau})}v_{i}(S_{i})\Big{]}$ $\geq\frac{1}{e^{2}k}\left(\frac{e-1}{e}\right)^{2}{\mathbb{E}}_{V}\Big{[}\sum_{e_{i}\in M^{V}\setminus\delta(H^{\tau})}\mathds{1}_{\{\chi_{i}\}}\mathds{1}_{\{\xi_{i}\}}\cdot\overline{w}_{i}\Big{]}$
$\geq\frac{1}{e^{3}k^{2}}\left(\frac{e-1}{e}\right)^{2}{\mathbb{E}}_{V}\Big{[}\sum_{e_{i}\in M^{V}\setminus\delta(H^{\tau})}\mathds{1}_{\{\chi_{i}\}}\cdot\overline{w}_{i}\Big{]}.$
Then, by adding contribution from buyers $i\in\delta(H^{\tau})$, we have

### v2

@@ 3.3.2 Reducing Game of Googol to $(\frac{1}{2},p)$-sample MRF
Our algorithm will partition the GoG instance with sample/real sets $(S,R)$ into two subinstances $(S_{1},R_{1})$ and $(S_{2},R_{2})$, and run the $(\frac{1}{2},p)$-sample algorithm on each subinstance. To create this partition, we will flip independent coins for each $i$. If coin $i$ is heads, we add $s_{i}$ to $S_{1}$ and $r_{i}$ to $R_{2}$. If tails, we add $s_{i}$ to $S_{2}$ and $r_{i}$ to $R_{1}$. This ensures $|S_{1}\cup R_{1}|=|S_{2}\cup R_{2}|=n$, so each subinstance is a valid input to the $(\frac{1}{2},p)$-sample algorithm. It now only remains to bound the performance of this algorithm.
Formally, let $\mathcal{I}$ be an instance of the Game of Googol model, and denote by ${\rm OPT}_{GoG}(\mathcal{I})$ and ${\rm ALG}_{GoG}(\mathcal{I})$ our final algorithm’s and the optimal solution cost for $\mathcal{I}$ respectively. Denote by $v^{1}_{1},\dots,v^{1}_{n}$ and $v^{-1}_{1},\dots,v^{-1}_{n}$ the values in $\mathcal{I}$, and by $\bm{\sigma}\sim\mathcal{D}_{\bm{\sigma}}$ the MRF where $\bm{\sigma}\in\{-1,1\}^{n}$. This defines a sample set $S=(v^{\sigma_{1}}_{1},\dots,v^{\sigma_{n}}_{n})$ and a real set $R=(v^{-\sigma_{1}}_{1},\dots,v^{-\sigma_{n}}_{n})$ in the instance $\mathcal{I}$.
To reduce to the $(\frac{1}{2},p)$-sample MRF setting, we will split the instance $(S,R)$ into two subinstances $(S_{1},R_{1})$ and $(S_{2},R_{2})$ as follows. First, we sample $\bm{\tilde{\sigma}}\sim\mathrm{Unif}\{-1,1\}^{n}$. Then, we define
$S_{1}$ $:=\{s_{i}\in S:\tilde{\sigma}_{i}=1\}$
$R_{1}$ $:=\{r_{i}\in R:\tilde{\sigma}_{i}=-1\}$
$S_{2}$ $:=\{s_{i}\in S:\tilde{\sigma}_{i}=-1\}$
$R_{2}$ $:=\{r_{i}\in R:\tilde{\sigma}_{i}=1\}$
Using these sets, we construct two instances: $\mathcal{I}_{1}=(S_{1},R_{1})$ and $\mathcal{I}_{2}=(S_{2},R_{2})$, where for $i\in\{1,2\}$, we have $S_{i}$ represent the sample set and $R_{i}$ represent the real set.
Starting with $\mathcal{I}_{1}$, we seek to show that the instance $(S_{1},R_{1})$ has a valid $(\frac{1}{2},p)$-sample MRF distribution conditional on any realization of $S_{1}\cup R_{1}=\{v_{i}^{\tilde{\sigma}_{i}\sigma_{i}}:i\in[n]\}$. Notice that for any value $v_{i}\in S_{1}\cup R_{1}$, we have:
$\Pr[v_{i}\in S_{1}\mid S_{1}\cup R_{1}]=\Pr[\sigma_{i}=1\mid(\tilde{\sigma}_{j}\sigma_{j})_{j\in[n]}]=\Pr[\sigma_{i}=1]=\frac{1}{2}$
by the definition of Game of Googol MRF and the independence of $(\tilde{\sigma}_{j}\sigma_{j})_{j\in[n]}$ and $\bm{\sigma}$, and
$\Pr[v_{i}\in S_{1}\mid S_{1}\setminus\{v_{i}\},~S_{1}\cup R_{1}]=\Pr[\sigma_{i}=1\mid\sigma_{-i},~(\tilde{\sigma}_{j}\sigma_{j})_{j\in[n]}]=\Pr[\sigma_{i}=1\mid\sigma_{-i}]\geq\frac{1}{2}\exp\left(-4\Delta\right)$
by Fact 2.3 and the aforementioned independence.
The equivalent property also holds for $\mathcal{I}_{2}$, so both $\mathcal{I}_{1}$ and $\mathcal{I}_{2}$ are (random) instances for the $(\frac{1}{2},p)$-sample MRF for $p=\frac{1}{2}\exp\left(-4\Delta\right)$. Denote by ${\rm ALG}_{MRF}$ the cost of the $\alpha(p)$-approximation algorithm for the $(\frac{1}{2},p)$-sample MRF model (we omit the $p$ for brevity) and denote by ${\rm OPT}_{MRF}$ the optimal cost.
@@
$\leq\alpha{\rm OPT}_{MRF}(\mathcal{I}_{1})+\alpha{\rm OPT}_{MRF}(\mathcal{I}_{2})$ $\alpha\text{-approx guarantee}$
$\leq\alpha{\rm OPT}_{MRF}(\mathcal{I}_{1}\cup\mathcal{I}_{2})+\alpha{\rm OPT}_{MRF}(\mathcal{I}_{1}\cup\mathcal{I}_{2})$ Monotonicity, (Assumption 3.1, (b))
$=2\alpha{\rm OPT}_{GoG}(\mathcal{I})$ $\text{Definition of }\mathcal{I}_{1},\mathcal{I}_{2}.$
@@ 4 Online Matchings and Combinatorial Auctions
For some distribution $V\sim\mathcal{D}$ and balanced prices $p^{V}_{1},\dots,p^{V}_{m}$, suppose there exist two posted-pricing schemes obtaining expected welfare ${\rm ALG}_{1}$ and ${\rm ALG}_{2}$, respectively, such that
@@
First, notice that Lemma 4.4 ensures we can always satisfy (1) for any distribution $\mathcal{D}$ that admits $(\alpha,\beta)$-balanced prices. For XOS valuations, Lemma 4.5 gives us $(\alpha,\beta)=(1,1)$-balanced prices that also have a pricing schemes satisfying requirement (2) of Lemma 4.3 with $\gamma=O(\Delta)$ and $\epsilon=1/e$. Similarly, for the hypergraph matching case, Lemma 4.5 gives us $(\alpha,\beta)=(1,k)$-balanced prices that also have a pricing scheme satisfying (2) with $\gamma=O(k^{2}(\Delta+\log k))$ and $\epsilon=\frac{1}{ek}$ which implies the $O((\Delta+\log k)k^{2})$-approximation of the theorem. ∎
@@ 4.1 Tail Contribution
Now, using our MRF conditioning lemma (2.3) and the fact that $R^{V}_{i}$ depends only on $v_{1},\dots,v_{i-1}$, we have ${\mathbb{E}}_{V^{\prime}}[\mathds{1}_{\{j\in O^{V^{\prime}}_{i}\setminus R^{V}_{i}\}}\cdot p^{V^{\prime}}_{j}]\geq e^{-4\Delta}{\mathbb{E}}_{V^{\prime}}[\mathds{1}_{\{j\in O^{V^{\prime}}_{i}\setminus R^{V}_{i}\}}\cdot p^{V^{\prime}}_{j}\mid v_{i}^{\prime}=v_{i}]$. Applying this to the above bound and rewriting $V^{\prime}$ as $\hat{V}$ to reflect the conditioning, we have
@@ 4.2.1 XOS Valuations
where each $\mathcal{A}_{i}$ is a subset of $\mathbb{R}_{\geq 0}^{m}$. We show, in Lemma 4.6, how to construct $(1,1)$-balanced prices that also obtain the welfare guarantee of Lemma 4.3.
@@ 4.2.2 Hypergraph Matching
where $M^{V}$ is the max weight hypergraph matching over $E=\{e_{i}:i\in[n]\}$ with weights $w_{e_{i}}=w_{i}$. In our main lemma we give $(1,k)$-balanced prices that also satisfy the requirements of Lemma 4.3, yielding the $O(k^{2}(\Delta+\log k))$ competitive ratio of Theorem 4.1.
@@
To obtain our prices, we first pick $\tau\sim Unif[0,4\Delta+\log k+2]$. Now, for each item $j$, let $\ell_{j}\in\mathbb{Z}$ be the unique integer such that $e^{\tau+\ell_{j}}\in[\frac{1}{e^{2}k}b_{j},~e^{4\Delta}b_{j})$ for $b_{j}:={\mathbb{E}}_{V}\left[p^{V}_{j}\right]$. Additionally, we sample $X_{\ell}\sim Ber(\frac{1}{k})$ independently for each $\ell\in\mathbb{Z}$ ; this variable indicates whether the item’s price will be set to high or low.
@@
$p_{j}:=\begin{cases}\frac{1}{e}\cdot e^{4\Delta}b_{j}&X_{\ell_{j}}=0\text{ (``high price''}),\\ \frac{1}{e}\cdot e^{\tau+\ell_{j}}&X_{\ell_{j}}=1\text{ (``low price'')}.\end{cases}$
First, we show that items which sell frequently at a high price achieve a high enough revenue to account for all their neighboring edges. Let $H^{t}:=\left\{j:\Pr[j\text{ sells}\mid X_{\ell_{j}}=0,~\tau=t]\geq\frac{1}{k^{2}}e^{-4\Delta-2}\right\}$ and denote by $\delta(H^{t})$ the neighboring edges of the items in the set $H^{t}$.
@@
For fixed $t$, we have ${\mathbb{E}}\big[\sum_{j\in H^{t}}\mathds{1}_{\{j\text{ sells}\}}\cdot p_{j}\mid\tau=t\big]\geq\frac{1}{2e^{3}k^{2}}\sum_{\begin{subarray}{c}e_{i}\in\delta(H^{t})\end{subarray}}\mathds{1}_{\{e_{i}\in M^{V}\}}\cdot w_{i}$
Proof of 4.8.
For $j\in H^{t}$, we have
${\mathbb{E}}[\mathds{1}_{\{j\text{ sells}\}}\cdot p_{j}\mid\tau=t]$ $\geq{\mathbb{E}}\left[\mathds{1}_{\{j\text{ sells}\}}\cdot e^{4\Delta-1}b_{j}\mid X_{\ell_{j}}=0,~\tau=t\right]\Pr[X_{\ell_{j}}=0\mid\tau=t]$
@@
Using $b_{j}={\mathbb{E}}\left[\sum_{e_{i}\in\delta(j)}\mathds{1}_{\{e_{i}\in M^{V}\}}w_{i}\right]$ and summing over $j\in H^{t}$, we get our claim. ∎
Hence, it suffices to focus on the contribution from buyers $i$ such that $e_{i}\not\in\delta(H^{\tau})$, as taking expectation over $\tau$ an applying 4.8 “covers” the value for $\delta(H^{\tau})$. For such a buyer $i$ with $e_{i}\not\in\delta(H^{\tau})$, define the $\xi_{i}$ to be the event that
@@
In other words, $\xi_{i}$ is the event that the items $j\in e_{i}$ with large $b_{j}$ are all “low-priced,” while all other items are “high-priced.” Let $B^{i}_{low}=\{j\in e_{i}:X_{\ell_{j}}=1\}$ and $B^{i}_{high}=\{j\in e_{i}:X_{\ell_{j}}=0\}$ denote these low and high priced items respectively. Notice that even in the event $\xi_{i}$, all items in $e_{i}$ have price at most $\frac{1}{e}e^{\tau+\ell^{i}_{\max}}$. Additionally, we note that $\xi_{i}$ is independent of the valuation $V$, and
@@
$\geq{\mathbb{E}}_{V}\Bigg[\mathds{1}_{\{\xi_{i}\}}\mathds{1}_{\{e_{i}\subseteq R^{V}_{i}\}}\cdot\Big(w_{i}-\frac{1}{e}e^{\tau+\ell^{i}_{\max}}\Big)^{+}\Bigg]$
@@
Here, $\chi_{i}$ denotes the event that $\frac{e^{\tau+\ell^{i}_{\max}}}{\overline{w}_{i}}\in[\frac{1}{e},1]$. We show the following claim on the term that corresponds to the high-priced items.
@@
${\mathbb{E}}\Big[\sum_{j\in[m]}\mathds{1}_{\{j\text{ sold}\}}\cdot p_{j}\Big]$ $\geq{\mathbb{E}}\Big[\sum_{j}\mathds{1}_{\{X_{\ell_{j}}=1\}}\mathds{1}_{\{j\text{ sold}\}}\cdot\frac{1}{e}e^{\tau+\ell_{j}}\Big]$
$={\mathbb{E}}\Big[\sum_{e_{i}\in M^{V}}\sum_{j\in e_{i}}\mathds{1}_{\{X_{\ell_{j}}=1\}}\mathds{1}_{\{j\text{ sold}\}}\cdot\frac{1}{e}e^{\tau+\ell_{j}}\Big]$
$={\mathbb{E}}\Big[\sum_{e_{i}\in M^{V}}\sum_{j\in B^{i}_{low}}\mathds{1}_{\{j\text{ sold}\}}\cdot\frac{1}{e}e^{\tau+\ell_{j}}\Big]$
$\geq{\mathbb{E}}\Big[\sum_{e_{i}\in M^{V}}\sum_{j\in B^{i}_{low}}\mathds{1}_{\{\xi_{i}\}}\mathds{1}_{\{j\text{ sold}\}}\cdot\frac{1}{e}e^{\tau+\ell^{i}_{\max}}\Big]$
$\geq{\mathbb{E}}\Big[\sum_{e_{i}\in M^{V}}\mathds{1}_{\{\xi_{i}\}}\mathds{1}_{\{B^{i}_{low}\not\subseteq R^{V}_{i}\}}\cdot\frac{1}{e}e^{\tau+\ell^{i}_{\max}}\Big]$
$\geq{\mathbb{E}}\Big[\sum_{e_{i}\in M^{V}}\mathds{1}_{\{\chi_{i}\}}\mathds{1}_{\{\xi_{i}\}}\mathds{1}_{\{B^{i}_{low}\not\subseteq R^{V}_{i}\}}\cdot\frac{1}{e^{2}}\overline{w}_{i}\Big].$
@@
${\mathbb{E}}\Big[\sum_{i\not\in\delta(H^{\tau})}v_{i}(S_{i})\Big]$ $\geq\frac{1}{e^{2}}\left(\frac{e-1}{e}\right)^{2}{\mathbb{E}}_{V}\Big[\sum_{e_{i}\in M^{V}\setminus\delta(H^{\tau})}\mathds{1}_{\{\chi_{i}\}}\mathds{1}_{\{\xi_{i}\}}\cdot\overline{w}_{i}\Big]$
$\geq\frac{1}{e^{3}k}\left(\frac{e-1}{e}\right)^{2}{\mathbb{E}}_{V}\Big[\sum_{e_{i}\in M^{V}\setminus\delta(H^{\tau})}\mathds{1}_{\{\chi_{i}\}}\cdot\overline{w}_{i}\Big].$
Then, by adding contribution from buyers $i\in\delta(H^{\tau})$ from 4.8, we have